# Inferring Outcome Means of Exponential Family Distributions Estimated by Deep Neural Networks

### v2

@@ A.1 Preliminaries
$\quad+\frac{r}{n}\sum_{k=1}^{r-1}\bigg{(}\frac{{{r-1}\choose{k}}}{{{n-1}\choose{k}}}-\frac{{{r}\choose{k+1}}}{{{n}\choose{k+1}}}\bigg{)}\sum_{2\leq i_{1}<...<i_{k}\leq n}T_{k+1}(\mathbf{z}_{1},\mathbf{z}_{i_{1}},...,\mathbf{z}_{i_{k}})$
$\quad+\frac{r}{n}\sum_{k=2}^{r}\bigg{(}\frac{{{r-1}\choose{k}}}{{{n-1}\choose{k}}}-\frac{{{r}\choose{k}}}{{{n}\choose{k}}}\bigg{)}\sum_{2\leq i_{1}<...<i_{k}\leq n}T_{k}(\mathbf{z}_{i_{1}},...,\mathbf{z}_{i_{k}}),$
and $M_{j}$ the same way with $1$ be replaced by $j$, and get that
$\widehat{\sigma}_{*}^{2}\sim\frac{n(n-1)}{(n-r)^{2}}\sum_{i=1}^{n}M_{i}^{2},$
similar to the proof of Theorem 3.7 which relies on the Hoeffding decomposition on $\widehat{f}^{b_{j}}(\mathbf{x}_{*})$ and $\widehat{f}^{B}(\mathbf{x}_{*})$. Therefore, denoting by
$\widehat{\sigma}_{1}^{2}=\frac{n(n-1)}{(n-r)^{2}}\sum_{i=1}^{n}M_{i}^{2},$
@@
$\mathbb{E}A_{i}\cdot T_{1}(\mathbf{z}_{i})=\mathbb{E}\bigg{(}\frac{r}{n}-\frac{r^{2}}{n^{2}}\bigg{)}T_{1}^{2}(\mathbf{z}_{1})=\bigg{(}\frac{r}{n}-\frac{r^{2}}{n^{2}}\bigg{)}\xi_{1,r}(\mathbf{x}_{*}).$ (A.7)
@@
$\mathbb{E}(M_{i}-A_{i})^{2}$ $=\frac{r^{2}}{n^{2}}\sum_{k=1}^{r-1}{{n-1}\choose{k}}\bigg{(}\frac{{{r-1}\choose{k}}}{{{n-1}\choose{k}}}-\frac{{{r}\choose{k+1}}}{{{n}\choose{k+1}}}\bigg{)}^{2}V_{k+1}+\frac{r^{2}}{n^{2}}\sum_{k=2}^{r}{{n-1}\choose{k}}\bigg{(}\frac{{{r-1}\choose{k}}}{{{n-1}\choose{k}}}-\frac{{{r}\choose{k}}}{{{n}\choose{k}}}\bigg{)}^{2}V_{k}$
$\precsim\frac{r^{2}(n-1)}{n^{2}}\bigg{(}\frac{{r-1}}{{n-1}}-\bigg{(}\frac{{\binom{r}{2}}}{{\binom{n}{2}}}\bigg{)}\bigg{)}^{2}\binom{r}{2}^{-1}\operatorname{{\rm Var}}[\widehat{f}^{b}(\mathbf{x}_{*})]$
$\precsim\frac{2r^{2}\operatorname{{\rm Var}}[\widehat{f}^{b}(\mathbf{x}_{*})]}{n^{3}}.$ (A.9)
Here, the first inequality is by the fact that summation in the first line is maximized in the second-order terms and $\sum_{k=1}^{r}{r\choose{k}}V_{k}=\operatorname{{\rm Var}}(\widehat{f}^{b}(\mathbf{x}_{*}))\leq C<+\infty$, and the second inequality is by simple calculation. We hence have
@@
$\qquad=\frac{n^{3}(n-1)}{r^{2}(n-r)^{2}}\bigg{(}\frac{r^{2}(n-r)^{2}}{n^{3}(n-1)}-\frac{2r^{2}(n-r)^{2}}{n^{4}}+\frac{r^{2}(n-r)^{2}}{n^{4}}\bigg{)}\to 0.$
Here, the last equality comes from (A.7) and (A.8). For any sequence of random variable $X_{n}>0$, $\mathbb{E}X_{n}\to 0$ implies $X_{n}~{}{\stackrel{{\scriptstyle p}}{{\rightarrow}}}~{}0$ by the Markov inequality. Therefore, we conclude that
$\frac{n}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}\cdot\frac{n(n-1)}{(n-r)^{2}}\sum_{i=1}^{n}\bigg{(}A_{i}-\frac{r(n-r)}{n^{2}}T_{1}(\mathbf{z}_{i})\bigg{)}^{2}~{}{\stackrel{{\scriptstyle p}}{{\rightarrow}}}~{}0.$ (A.10)
@@
$\frac{n}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}\cdot\frac{n(n-1)}{(n-r)^{2}}\sum_{i=1}^{n}\frac{r^{2}(n-r)^{2}}{n^{4}}T_{1}^{2}(\mathbf{z}_{i})~{}{\stackrel{{\scriptstyle p}}{{\rightarrow}}}~{}1.$ (A.11)
@@
Here, we use the fact $\mathbb{E}(\sum_{i^{\prime}\neq i}^{n}T_{1}(\mathbf{z}_{i^{\prime}})T_{1}(\mathbf{z}_{i}))^{2}=n(n-1)\xi_{1,r}^{2}(\mathbf{x}_{*})$ which follows from a straightforward calculation. This implies that the interaction term converges to 00 in probability. With these two terms, we have
$\frac{n}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}\cdot\frac{n(n-1)}{(n-r)^{2}}\sum_{i=1}^{n}\frac{r(n-r)}{n^{2}}A_{i}\cdot T_{1}(\mathbf{z}_{i})~{}{\stackrel{{\scriptstyle p}}{{\rightarrow}}}~{}1.$ (A.12)
Combing (A.10) and (A.11) with (A.12), we have
@@
Moreover, it follows from (A.9) that
@@
Appendix B Proofs in Appendix A
We provide additional technical proofs for the conclusions in Appendix A. We first present the proof of Lemma A.4. Then we show several additional lemmas, which will be used to prove Proposition A.1.
Proof of Lemma A.4.

### v3

@@ C.1 Preliminaries
$\quad+\frac{r}{n}\sum_{k=1}^{r-1}\bigg(\frac{{\binom{r-1}{k}}}{{\binom{n-1}{k}}}-\frac{{\binom{r}{k+1}}}{{\binom{n}{k+1}}}\bigg)\sum_{2\leq i_{1}<...<i_{k}\leq n}T_{k+1}(\mathbf{z}_{1},\mathbf{z}_{i_{1}},...,\mathbf{z}_{i_{k}})$
$\quad+\frac{r}{n}\sum_{k=2}^{r}\bigg(\frac{{\binom{r-1}{k}}}{{\binom{n-1}{k}}}-\frac{{\binom{r}{k}}}{{\binom{n}{k}}}\bigg)\sum_{2\leq i_{1}<...<i_{k}\leq n}T_{k}(\mathbf{z}_{i_{1}},...,\mathbf{z}_{i_{k}}),$
and $M_{i}$ the same way with index $1$ of $\mathbf{z}$ be replaced by $i$ of $\mathbf{z}$. Let the whole data set $\mathcal{D}_{n}=\{\mathbf{z}_{1},...,\mathbf{z}_{n}\}$, and recall $Z_{b_{j}i}=(J_{b_{j}i}-J_{\cdot i})(\widehat{f}^{b_{j}}(\mathbf{x}_{*})-\widehat{f}^{B}(\mathbf{x}_{*}))$, the motivation of the definition of $M_{i}$ is given by the direct calculation that
$\mathbb{E}(Z_{b_{j}i}|\mathcal{D}_{n})=M_{i}.$
Here, the expectation takes over the randomness of subsampling. Define
$\widehat{\sigma}_{1}^{2}=\frac{n(n-1)}{(n-r)^{2}}\sum_{i=1}^{n}M_{i}^{2},\quad R_{B}=\widetilde{\sigma}^{2}-\widehat{\sigma}_{1}^{2}.$
We next prove two key conclusions, which will show the consistency of our results.
$\begin{aligned} &\frac{n(n-1)}{(n-r)^{2}}\frac{1}{B(B-1)}\sum_{i=1}^{n}\sum_{j=1}^{B}\bigl(Z_{b_{j}i}-\widehat{V}_{i}\bigr)^{2}-R_{B}=o_{p}\bigg(\frac{r^{2}\xi_{1,r}(\mathbf{x}_{*})}{n}\bigg),\\ &\sqrt{\frac{r^{2}\xi_{1,r}(\mathbf{x}_{*})}{n\widehat{\sigma}_{1}^{2}}}~{\stackrel{{\scriptstyle p}}{{\rightarrow}}}~1,\end{aligned}$ (C.7)
In the following of the proof, we denote by $\widehat{R}_{B}=\frac{n(n-1)}{(n-r)^{2}}\frac{1}{B(B-1)}\sum_{i=1}^{n}\sum_{j=1}^{B}\bigl(Z_{b_{j}i}-\widehat{V}_{i}\bigr)^{2}$. If (C.7) holds, we can see that
$\frac{n\widehat{\sigma}_{*}^{2}}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}=\frac{n(\widetilde{\sigma}^{2}-\widehat{R}_{B})}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}=\frac{n(R_{B}-\widehat{R}_{B})}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}+\frac{n(\widetilde{\sigma}^{2}-R_{B})}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}=o_{p}(1)+\frac{n\widehat{\sigma}_{1}^{2}}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}~{\stackrel{{\scriptstyle p}}{{\rightarrow}}}~1,$
which completes the proof of Theorem 3.8. It remains for us to prove (C.7).
We first investigate the term $\widehat{R}_{B}-R_{B}=\frac{n(n-1)}{(n-r)^{2}}\frac{1}{B(B-1)}\sum_{i=1}^{n}\sum_{j=1}^{B}\bigl(Z_{b_{j}i}-\widehat{V}_{i}\bigr)^{2}-R_{B}$ in (C.7). Recall $R_{B}=\widetilde{\sigma}^{2}-\widehat{\sigma}_{1}^{2}$, it is easy to see that
$R_{B}=\frac{n(n-1)}{(n-r)^{2}}\sum_{i=1}^{n}(\widehat{V}_{i}^{2}-M_{i}^{2}).$
Let $\varepsilon_{b_{j}i}=Z_{b_{j}i}-M_{i}$, and $\overline{\varepsilon}_{i}=\frac{1}{B}\sum_{j=1}^{B}\varepsilon_{b_{j}i}$, recall $\widehat{V}_{i}=\frac{\sum_{j=1}^{B}Z_{b_{j}i}}{B}$, we have that $\widehat{V}_{i}=M_{i}+\overline{\varepsilon}_{i}$ and $\widehat{V}_{i}^{2}-M_{i}^{2}=2M_{i}\overline{\varepsilon}_{i}+\overline{\varepsilon}_{i}^{2}$, which means
$R_{B}=\frac{n(n-1)}{(n-r)^{2}}\sum_{i=1}^{n}2M_{i}\overline{\varepsilon}_{i}+\overline{\varepsilon}_{i}^{2}.$
Further denote by $\widehat{\mathrm{Var}}_{i}=\frac{1}{B-1}\sum_{j=1}^{B}\bigl(Z_{b_{j}i}-\widehat{V}_{i}\bigr)^{2}$, we have
$\widehat{R}_{B}=\frac{n(n-1)}{(n-r)^{2}}\frac{1}{B}\sum_{i=1}^{n}\widehat{\mathrm{Var}}_{i}.$
We then combine the expression of $\widehat{R}_{B}$ and $R_{B}$ above and get that
$R_{B}-\widehat{R}_{B}=\frac{n(n-1)}{(n-r)^{2}}\bigg(\underbrace{\sum_{i=1}^{n}2M_{i}\overline{\varepsilon}_{i}}_{\Delta_{1}}+\underbrace{\sum_{i=1}^{n}\overline{\varepsilon}_{i}^{2}-B^{-1}\mathrm{Var}_{i}}_{\Delta_{2}}-\underbrace{\frac{1}{B}\sum_{i=1}^{n}\widehat{\mathrm{Var}}_{i}-\mathrm{Var}_{i}}_{\Delta_{3}}\bigg).$
Here, $\mathrm{Var}_{i}=\mathrm{Var}_{b}(Z_{b_{j}i})$ with variance condition on the dataset $\mathcal{D}_{n}$. We bound the three terms one by one. For the term $\Delta_{1}$, by the definition of $M_{i}$ and $\overline{\varepsilon_{i}}$, we have
$\Delta_{1}=\frac{2}{B}\sum_{j=1}^{B}\sum_{i=1}^{n}M_{i}\varepsilon_{b_{j}i}.$
Conditional on the dataset $\mathcal{D}_{n}$, $\sum_{i=1}^{n}M_{i}\varepsilon_{b_{j}i}$ is independent with different $b_{j}$. It is easy to see that $\mathbb{E}_{b}\Delta_{1}=0$, and we give the calculation of $\mathrm{Var}_{b}\Delta_{1}$. From the expression of $\Delta_{1}$, conditional on data set $\mathcal{D}_{n}$, $\varepsilon_{b_{j}i}$ are independent for different $b_{j}$, it is easy to see that
$\mathrm{Var}_{b}\Delta_{1}\asymp\frac{\mathrm{Var}_{b}\sum_{i=1}^{n}M_{i}\varepsilon_{b_{j}i}}{B}.$
It remains for us to investigate the term in the numerator to bound $\Delta_{1}$. We have
$\mathrm{Var}_{b}\sum_{i=1}^{n}M_{i}\varepsilon_{b_{j}i}=\sum_{i=1}^{n}M_{i}^{2}\mathrm{Var}_{b}\varepsilon_{b_{j}i}+\sum_{i\neq i^{\prime}}M_{i}M_{j}\operatorname{\rm Cov}_{b}(\varepsilon_{b_{j}i},\varepsilon_{b_{j}i^{\prime}}).$
From the definition of $M_{i}$, we can easily bound $M_{i}=O(r/n)$, for the covariance of $\varepsilon_{b_{j}i}$ and $\varepsilon_{b_{j}i^{\prime}}$, simple calculation of the subsampling randomness gives us that $\mathrm{Var}_{b}\varepsilon_{b_{j}i}=O(\frac{r}{n}(1-\frac{r}{n})=O(r/n)$, and $\operatorname{\rm Cov}_{b}(\varepsilon_{b_{j}i},\varepsilon_{b_{j}i^{\prime}})=O(|\frac{r(r-1)}{n(n-1)}-\frac{r^{2}}{n^{2}})|=O(\frac{r}{n^{2}})$, we hence have
$\mathrm{Var}_{b}\sum_{i=1}^{n}M_{i}\varepsilon_{b_{j}i}$ $=O\bigg(n\cdot\frac{r^{2}}{n^{2}}\cdot\frac{r}{n}+n^{2}\cdot\frac{r^{2}}{n^{2}}\cdot\frac{r}{n^{2}}\bigg)$
$=O\bigg(\frac{r^{3}}{n^{2}}\bigg).$
This indicates that $\mathrm{Var}_{b}\Delta_{1}=O(\frac{r^{3}}{Bn^{2}})$. Combining the fact that $\mathbb{E}_{b}\Delta_{1}=0$, we conclude that $|\Delta_{1}|=O_{p}(\frac{r^{3/2}}{n\sqrt{B}})$.
For the term $\Delta_{2}$, by the expression that $\overline{\varepsilon}_{i}^{2}=\frac{1}{B^{2}}\sum_{j=1}^{B}\varepsilon_{b_{j}i}^{2}+\frac{1}{B^{2}}\sum_{j\neq j^{\prime}}\varepsilon_{b_{j}i}\varepsilon_{b_{j^{\prime}}i}$, $\Delta_{2}$ can be further written as
$\Delta_{2}=\underbrace{\frac{1}{B^{2}}\sum_{i=1}^{n}\sum_{j=1}^{B}\big(\varepsilon_{b_{j}i}^{2}\big)-B^{-1}\mathrm{Var}_{i}}_{\Delta_{21}}+\underbrace{\frac{1}{B^{2}}\sum_{i=1}^{n}\sum_{j\neq j^{\prime}}\varepsilon_{b_{j}i}\varepsilon_{b_{j^{\prime}}i}}_{\Delta_{22}}.$
We have $\mathbb{E}_{b}\Delta_{21}=\mathbb{E}_{b}\Delta_{22}=0$. Simple calculation based on the randomness of subsampling gives $\mathrm{Var}_{b}(\Delta_{21})=O(n^{2}/B^{3})$, we have $|\Delta_{21}|=O_{p}(n/B^{3/2})$. As for $\Delta_{22}$, we consider $T_{b_{j},b_{j^{\prime}}}=\sum_{i=1}^{n}\varepsilon_{b_{j}i}\varepsilon_{b_{j^{\prime}}i}$, and we have $\mathbb{E}_{b}T_{b_{j},b_{j^{\prime}}}^{2}=O(r^{2}/n)$, and $\mathbb{E}_{b}T_{b_{j},b_{j^{\prime}}}\cdot T_{b^{\prime}_{j},b^{\prime}_{j^{\prime}}}=0$ as long as $b_{j},b_{j^{\prime}},b^{\prime}_{j},b^{\prime}_{j^{\prime}}$ has three or more distinct values by the conditional independence. We have
$\mathrm{Var}_{b}(\Delta_{22})\leq\frac{1}{B^{4}}\cdot\mathbb{E}_{b}\sum_{b_{j}\neq b_{j^{\prime}},b^{\prime}_{j}\neq b^{\prime}_{j^{\prime}}}T_{b_{j},b_{j^{\prime}}}\cdot T_{b^{\prime}_{j},b^{\prime}_{j^{\prime}}}=O\bigg(\frac{r^{2}}{nB^{2}}\bigg)$
We conclude that $|\Delta_{2}|=O_{p}(n/B^{3/2}+r/(B\sqrt{n}))$.
For the term $\Delta_{3}$, by the definition of $\widehat{\mathrm{Var}}_{i}=\frac{1}{B-1}\sum_{j=1}^{B}\bigl(Z_{b_{j}i}-\widehat{V}_{i}\bigr)^{2}$, it holds that
$\Delta_{3}$ $=\frac{1}{B(B-1)}\sum_{i=1}^{n}\sum_{j=1}^{B}\bigg((Z_{b_{j}i}-\widehat{V}_{i})^{2}-\frac{B-1}{B}\mathrm{Var}_{i}\bigg)$
$=\bigg(\frac{1}{B(B-1)}\sum_{i=1}^{n}\sum_{j=1}^{B}\bigg((Z_{b_{j}i}-M_{i})^{2}-\mathrm{Var}_{i}\bigg)\bigg)+O\bigg(\frac{n}{B^{2}}\bigg).$
Here, the first equality is by direct calculation and the second equality is by $\mathrm{Var}_{i}$ is bounded. Let , easy to bound that $\mathrm{Var}_{b}(T_{b_{j}})=O(n^{2})$, we have that
$\mathbb{E}_{b}\bigg\{\frac{1}{B(B-1)}\sum_{i=1}^{n}\sum_{j=1}^{B}\bigg((Z_{b_{j}i}-M_{i})^{2}-\mathrm{Var}_{i}\bigg)\bigg\}=0;$
$\mathrm{Var}_{b}\bigg\{\frac{1}{B(B-1)}\sum_{i=1}^{n}\sum_{j=1}^{B}\bigg((Z_{b_{j}i}-M_{i})^{2}-\mathrm{Var}_{i}\bigg)\bigg\}=O(n^{2}/B^{3}),$
we hence conclude that $|\Delta_{3}|=O_{p}(n/B^{3/2})$.
We collect all the error terms of $\Delta_{1}$, $\Delta_{2}$ and $\Delta_{3}$, and have
$\Delta_{1}=O_{p}\bigg(\frac{r^{3/2}}{n\sqrt{B}}\bigg),\quad\Delta_{2}=O_{p}(n/B^{3/2}+r/(B\sqrt{n})),\quad\Delta_{3}=O_{p}(n/B^{3/2}),$
easy from order comparison to see that if $\Delta_{1}=o_{p}\big(\frac{r^{2}\xi_{1,r}(\mathbf{x}_{*})}{n}\big)$, then $R_{B}-\widehat{R}_{B}=o_{p}\big(\frac{r^{2}\xi_{1,r}(\mathbf{x}_{*})}{n}\big)$. We have
$\frac{n\Delta_{1}}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}=\frac{1}{\sqrt{rB}\xi_{1,r}(\mathbf{x}_{*})}=o_{p}(1),$
where we use the Assumption 3.4 that $B\succsim n$ and $\varepsilon$ small enough. We hence conclude that $R_{B}-\widehat{R}_{B}=o_{p}\big(\frac{r^{2}\xi_{1,r}(\mathbf{x}_{*})}{n}\big)$.
@@
$\mathbb{E}A_{i}\cdot T_{1}(\mathbf{z}_{i})=\mathbb{E}\bigg(\frac{r}{n}-\frac{r^{2}}{n^{2}}\bigg)T_{1}^{2}(\mathbf{z}_{1})=\bigg(\frac{r}{n}-\frac{r^{2}}{n^{2}}\bigg)\xi_{1,r}(\mathbf{x}_{*}).$ (C.8)
@@
$\mathbb{E}(M_{i}-A_{i})^{2}$ $=\frac{r^{2}}{n^{2}}\sum_{k=1}^{r-1}{\binom{n-1}{k}}\bigg(\frac{{\binom{r-1}{k}}}{{\binom{n-1}{k}}}-\frac{{\binom{r}{k+1}}}{{\binom{n}{k+1}}}\bigg)^{2}V_{k+1}+\frac{r^{2}}{n^{2}}\sum_{k=2}^{r}{\binom{n-1}{k}}\bigg(\frac{{\binom{r-1}{k}}}{{\binom{n-1}{k}}}-\frac{{\binom{r}{k}}}{{\binom{n}{k}}}\bigg)^{2}V_{k}$
$\precsim\frac{r^{2}(n-1)}{n^{2}}\bigg(\frac{{r-1}}{{n-1}}-\bigg(\frac{{\binom{r}{2}}}{{\binom{n}{2}}}\bigg)\bigg)^{2}\binom{r}{2}^{-1}\mathrm{Var}[\widehat{f}^{b}(\mathbf{x}_{*})]$
$\precsim\frac{2r^{2}\mathrm{Var}[\widehat{f}^{b}(\mathbf{x}_{*})]}{n^{3}}.$ (C.10)
Here, the first inequality is by the fact that summation in the first line is maximized in the second-order terms and $\sum_{k=1}^{r}{\binom{r}{k}}V_{k}=\mathrm{Var}(\widehat{f}^{b}(\mathbf{x}_{*}))\leq C<+\infty$, and the second inequality is by simple calculation. We hence have
@@
$\qquad=\frac{n^{3}(n-1)}{r^{2}(n-r)^{2}}\bigg(\frac{r^{2}(n-r)^{2}}{n^{3}(n-1)}-\frac{2r^{2}(n-r)^{2}}{n^{4}}+\frac{r^{2}(n-r)^{2}}{n^{4}}\bigg)=\frac{1}{n}\to 0.$
Here, the last second equality comes from (C.8) and (C.9). For any sequence of random variable $X_{n}>0$, $\mathbb{E}X_{n}\to 0$ implies $X_{n}~{\stackrel{{\scriptstyle p}}{{\rightarrow}}}~0$ by the Markov inequality. Therefore, we conclude that
$\frac{n}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}\cdot\frac{n(n-1)}{(n-r)^{2}}\sum_{i=1}^{n}\bigg(A_{i}-\frac{r(n-r)}{n^{2}}T_{1}(\mathbf{z}_{i})\bigg)^{2}~{\stackrel{{\scriptstyle p}}{{\rightarrow}}}~0.$ (C.11)
@@
$\frac{n}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}\cdot\frac{n(n-1)}{(n-r)^{2}}\sum_{i=1}^{n}\frac{r^{2}(n-r)^{2}}{n^{4}}T_{1}^{2}(\mathbf{z}_{i})~{\stackrel{{\scriptstyle p}}{{\rightarrow}}}~1.$ (C.12)
@@
Here, we use the fact $\mathbb{E}(\sum_{i^{\prime}\neq i}^{n}T_{1}(\mathbf{z}_{i^{\prime}})T_{1}(\mathbf{z}_{i}))^{2}=n(n-1)\xi_{1,r}^{2}(\mathbf{x}_{*})$ which follows from a straightforward calculation. This implies that the interaction term converges to 0 in probability. With these two terms, we have
$\frac{n}{r^{2}\xi_{1,r}(\mathbf{x}_{*})}\cdot\frac{n(n-1)}{(n-r)^{2}}\sum_{i=1}^{n}\frac{r(n-r)}{n^{2}}A_{i}\cdot T_{1}(\mathbf{z}_{i})~{\stackrel{{\scriptstyle p}}{{\rightarrow}}}~1.$ (C.13)
Combing (C.11) and (C.12) with (C.13), we have
@@
Moreover, it follows from (C.10) that
@@
Appendix D Proofs in Section C
We provide additional technical proofs for the conclusions in Section C. We first present the proof of Lemma C.4. Then we show several additional lemmas, which will be used to prove Proposition C.1.
Proof of Lemma C.4.